OpenStudy (anonymous):

graph the equation by plotting three points x -4y=-18

OpenStudy (anonymous):

From this equation you can find infinite number of points. Do so find plugging in 0,1,-1 for x and looking for y

OpenStudy (anonymous):

i dont get this?

OpenStudy (anonymous):

x - 4y=18

OpenStudy (anonymous):

its supposed to be x - 4y=-18

OpenStudy (anonymous):

y intercept...y=18/4 x intercept...x=18 its a straight line with a gradient of 1/4 so its an increasing straight line

OpenStudy (anonymous):

what?....... y =18,4

OpenStudy (anonymous):

its supposed to be 3 points

OpenStudy (anonymous):

the only thing you can get from the graph is gradient and intercepts

OpenStudy (anonymous):

oh okhay i dont know what gradients mean

OpenStudy (anonymous):

to get 3 points do as follws: first put x=2 and u get 2 -4y = -18 -4y = -18 -2 y = -20/-4 y = 5 so first point is (x,y) = (2,5)

OpenStudy (anonymous):

Slope

OpenStudy (anonymous):

first put x= -2 and u get -2 -4y = -18 -4y = -18 +2 y = -16/-4 y = 4 so second point is (x,y) = (-2,4)

OpenStudy (anonymous):

okhay what are the next 2 points?

OpenStudy (anonymous):

lol, girlie, you dont know what a gradient is?, OMG, thats suprising

OpenStudy (anonymous):

wow it surprised me too.

OpenStudy (anonymous):

finallyput x= -6 and u get -6 -4y = -18 -4y = -18 + 6 y = -12/-4 y = 3 so third point is (x,y) = (-6,3) Now plot these three points and join them with a line to get the required graph of the equation

OpenStudy (anonymous):

Gradient is called Slope is U.S

OpenStudy (anonymous):

@minah got it??????

OpenStudy (anonymous):

ummm im from america ...... yes harkirat i got it you deserve a medal thanks sweetie

OpenStudy (anonymous):

whenever they ask you to plot by taking three points follows this procedure...

OpenStudy (anonymous):

@HARKIRAT okhay but i really suck at it =)

OpenStudy (anonymous):

O SORRY GIRL YOU NEED MORE PRACTICE BECAUSE IM ALSO FROM THE USA.

OpenStudy (anonymous):

YEA YOU ARE SORRY , BUT UMM JUST BECAUSE I DONT KNOW WHAT THAT MEANS DON'T MEAN NOTHING OKHAY? EINSTEIN

OpenStudy (anonymous):

ok, stop with all caps

OpenStudy (anonymous):

ok goodluck in your math practice pray be4 you study, trust me, it works,God bless!!!

OpenStudy (anonymous):

thanks guys/girls for the help =) everyone!!

OpenStudy (anonymous):

@einstien ---- All fingers are not alike and no two minds are alike. Can u sing like MJ or dance as good as MJ Can u play tennis like Roger or Nadal?? Each of us has our own pace of learning as long as we keep trying and improving, it is okay..... PS : You seem to have problems with spellings!!! Its Einstein and not eistein !!

OpenStudy (anonymous):

wow nicely said...

OpenStudy (anonymous):

Do I get a medal for this too !!! :D:D:D

OpenStudy (anonymous):

yeahhh i already gave you one i dont know how to give you another ... but you could help me more when i ask for questions for more medals

OpenStudy (anonymous):

just joking......... if i am online i will definitely help.....

OpenStudy (anonymous):

you know great mathematicians and scientists like me dont care about the grammar but care about the figures instead, sorry.

OpenStudy (anonymous):

ok i wil give every one a medal

OpenStudy (anonymous):

I think u hv a long way before u can b called "great mathematician and scientist". win a Nobel Prize for something that u hv done and i will bow before u !! Being able to do successfully something that others have developed is no genius !!!

OpenStudy (anonymous):

my friend i hate essays go and write a novel for my granny to read

OpenStudy (anonymous):

Oh so others have to take of your granny's needs !! don't u care for her??

OpenStudy (anonymous):

omg GUYS please stop!!!!!!! your both smart people...... don't argue over dumb things =)

OpenStudy (anonymous):

lol dude cumon. lets stop.

OpenStudy (anonymous):

minah, he has chosen the names of great but very humble people but seems to be too full of himself. it is not good to look down upon others.....that's all

OpenStudy (anonymous):

Ok Peace !!!

OpenStudy (anonymous):

thanks help me again =)